OpenStudy (anonymous):

Find the slope of line passing through the points (4,8) and (6, -2).

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

plz fan and give me best response

OpenStudy (texaschic101):

let me show you HOW to do this. Use the slope formula : slope = (y2 - y1) / (x2 - x1) (4,8)...x1 = 4 and y1 = 8 (6,-2)...x2 = 6 and y2 = -2 now we sub slope = (-2 - 8) / (6 - 4) slope = -10/2 slope = -5

OpenStudy (anonymous):

thanks for the insight,